OpenStudy (anonymous):

The perimeter of a rectangle is 48 m. If the width were doubled and the length were increased by 24 m, the perimeter would be 112 m. What is the length of the original rectangle?

OpenStudy (anonymous):

let w be width and l be length 2w+2l=48 2l=48-2w 4w+2(l+24)=112 4w+2l+48=112 4w+48-2w=64 2w=16 w=8 2l=48-16 =32 l=16

OpenStudy (anonymous):

let l=length w=width 2(l+w)=48 l+w=24 l=-w+24 (1) 2(l+w)=112 2[(l+24)+2w]=112 2l+48+4w=112 (2) sustitute the value of length 2(-w+24)+48+4w=112 -2w+48+48+4w =112 2w+96 =112 2w =112-96 2w =16 w= 8 l=16